OpenStudy (anonymous):

Which is a factored form of 27x^3+64

OpenStudy (raden):

27x^3 = (3x)^3 and 64 = 4^3 so, 27x^3+64 = (3x)^3 + 4^3 now you can use this formula : a^3 + b^3 = (a+b)(a^2 - ab + b^3) in this case a = 3x and b = 4